OpenStudy (anonymous):

PLease help me with this i am having soo much truble with it A family is having a pool built in their backyard. If their yard is rectangular and measures 7x by 6x and the pool is circular with a radius of 2x how much of the yard will be left over after the pool is built? 38x2 42x2 – 4x2 42x2 – 2x2 38x2

OpenStudy (anonymous):

hmm... looks like the answer is neither of those choices.... the answer should be in terms of pi....

OpenStudy (anonymous):

those are the answer choices..

OpenStudy (anonymous):

ok... sorry i can't help...

OpenStudy (anonymous):

this lesson is over Polynomials and Factoring

OpenStudy (anonymous):

ok its fine it confused the crap outta me lol

OpenStudy (anonymous):

just to close it out, Area left over = Area of Rectangle - Area of Circle Area of rectangle = (7x)(6x) = \(\ 42x^2 \) Area of circle = \(\ \pi \cdot r^2=\pi (2x)^2=4\pi x^2 \) so Area left over = \(\large 42x^2-4\pi x^2 \)

OpenStudy (anonymous):

thx lol i thought you couldnt help?

OpenStudy (anonymous):

i couldn't help with the answer choices but i can actually help in solving the problem...

OpenStudy (anonymous):

Can you help with these too? A sports team is building a neW stadium on a rectangular lot of land. If the lot measures 6x by 10x and the sports field will be 1x by 4x, how much of the lot will be left over to build bleachers on? 56x2 64x2 30x2 60x2 Simplify the product using the distributive property. 7. (5h – 3)(3h + 7) 15h2 – 44h + 21 15h2 – 26h – 21 15h2 + 44h + 21 15h2 + 26h – 21

OpenStudy (anonymous):

for the first one, just subtract.... Whole Field = (6x)(10x) = ??? Sports Field = (1x)(4x) = ??? Bleachers = Whole Field - Sports Field = ??? for #7, do you know FOIL ???

OpenStudy (anonymous):

no lol whats foil?

OpenStudy (anonymous):

and for the first one idk how to subtact those.

OpenStudy (anonymous):

FOIL : First, Outside, Inside, Last